# Quantitative evaluation of methods to analyze motion changes in single-particle experiments

### v2

@@ Methods
  • Field of view (FOV) — A region of the sample where the recording takes place. Particles within the same FOV can undergo interactions among themselves and/or with the environment.
  • Video (Track 1 only) — Videos corresponding to each FOV.
  • Trajectory (Track 2 only) — Trajectory corresponding to the motion of an individual particle.
@@
For both tracks, all the particles used in the simulations are provided/visualized (i.e., full labeling conditions). The effect of blinking or photobleaching is not taken into account.
Figure 2: Physical models of interaction and structure of the simulated datasets. a, Examples of 2-dimensional trajectories undergoing interactions inducing changes in their motion. From left to right: single-state model (SSM) without changes of diffusion; multi-state model (MSM) with time-dependent changes between different diffusive states (red and blue); dimerization model (DIM) where a particle (red) selectively interacts with another particle (gray) and the two transiently co-diffuse with a different motion (blue trajectory); transient-confinement model (TCM) where a particle diffuses inside (blue) and outside (red) compartments with osmotic boundaries (gray area); quenched-trap model (QTM) where a particle is transiently immobilized (blue) at specific loci through interactions with static features of the environment (gray areas). b An experiment (left panel) consists of simulations performed according to one of the models of interactions described in a (here shown a TCM experiment), with a set of parameters describing the dynamic interplay of the particles and/or the environment. From the same experiment, several fields of view (FOVs) are selected. Particles within the same FOV (right panel) diffuse and undergo interactions among themselves and/or with the environment (gray areas) that affect their trajectories. c, Time traces of the coordinates of exemplary trajectories from the experiment depicted in b displaying changes of diffusion properties at specific times (dashed vertical lines). For the challenge, the motion analysis can be either performed directly from the video recording of the FOV (Track 1), or from detected trajectories linking the coordinates of individual particles at different times (Track 2).
@@
In each track, participants can compete in two different tasks, as typically done in the analysis of experimental data:
@@
  • Ensemble Task — Ensemble-level predictions providing for each experimental condition the model used to simulate the experiment, the number of states, and the fraction of time spent in each state. For each identified state, participants should determine the mean and standard deviation of the distribution of the generalized diffusion coefficients $K$, and the mean and standard deviation of the distribution of the anomalous diffusion exponent $\alpha$ corresponding to the underlying motion.
@@
  • Single-trajectory Task — Trajectory-level predictions providing for each trajectory a list of $M$ inner CPs delimiting $M+1$ segments with different dynamic behavior. For each segment, participants should identify the generalized diffusion coefficient $K$, the anomalous diffusion exponent $\alpha$ corresponding to the underlying motion, and an identifier of the kind of constraint imposed by the environment ($0=$ immobile, $1=$ confined, $2=$ free (unconstrained), $3=$ directed ($\alpha\geq 1.9$)). For Track 1, predictions must be provided for a subset of particles (in the following, we will refer to them as VIP, very important particles) identified through a label map of the first frame of the movie. For Track 2, predictions must be provided for all trajectories in the FOV.
@@
For each task, several metrics will be evaluated (see section Scoring and evaluations). Participants can provide partial submissions, e.g., including predictions for a limited subset of experiments or for specific parameters. For ranking purposes of the challenge, missing predictions will be scored with the worst possible value of the corresponding metric.
Competition timeline
[…]
The Development Phase (December 1, 2023, to April 30, 2024) is intended for the participants to set up their methods, test them, and familiarize themselves with the datasets and the scoring platform. An unlabeled dataset will be available and the public leaderboard will show scores obtained on this dataset. The code to generate further data will also be available.
The Validation Phase (May 1, 2024, to June 30, 2024) will be a test of the actual final challenge. A new dataset will be provided and the public leaderboard will show the scores for each task.
The Challenge Phase (July 1, 2024, to July 15, 2024) will be the final stage of the competition. Only teams in the top 10 ranking of each task will be allowed to participate in this phase. A new dataset will be provided and the number of submissions will be limited to 1 per day. Results obtained by other participants will not be publicly disclosed, as the leaderboard will be made publicly available only after the deadline.
The analysis of the final results will be performed from July to October 2024. If Stage 1 review is successful, the manuscript will be submitted by the end of October 2024. Members of the teams in the top-5 ranking of each task will be invited to contribute to the article as authors.
@@
The results of the challenge will be discussed with the participants and other experts from the field during the $2^{\rm nd}$ Anomalous Diffusion Workshop that will be held in December 2024.
Dataset structure
[…]
Figure 3: Structure of the dataset. Datasets for each track are contained in the respective directories, namely, Track 1 and Track 2. These include a folder for each experiment labeled with a sequential number (EXP_[exp number]). Each experiment folder contains a list of 30 files labeled with a sequential number (FOV_[fov number]) associated with different non-overlapping FOVs. For Track 1, the files consist of 200-frame videos and an additional map corresponding to the segmentation of VIP particles at the first frame. For Track 2, the files contain a table whose columns correspond to the trajectory index, time step, $x$-coordinate, and $y$-coordinate.
Figure 4: Examples of challenge datasets. a, b, Frames of exemplary videos (FOVs) provided for Track 1 corresponding to the different diffusion models and reproducing low (a) and high number density (b). c, d, The upper rows show plots of the trajectories for exemplary FOVs provided for Track 2 corresponding to the different diffusion models and different sets of parameters exhibiting evident (c) or subtle changes (d) of diffusion, i.e., data in c reproduce conditions were diffusion changes are easier to identify as compared to d. For example, for the SSM, all trajectories in c have the same generalized diffusion coefficient, whereas in d they have a broad distribution of $K$. For the MSM, trajectories in c undergo large changes of diffusivity and show long dwell times in each state, as compared to d. For the DIM, trajectories in c undergo large changes of anomalous diffusion exponent and show long dwell times in each state, as compared to d. For the TCM, trajectories cannot leave the confinement zone once they get inside c, whereas in d the compartments have a finite transmittance. For the QTM, the dwell times in the traps are much longer in c than in d. The lower rows show time traces of the $x$-coordinate of representative trajectories. The different colors indicate different diffusive states.
For Track 1 (3(a)), the coordinates of particles in the same FOV are used to generate 200-frame videos as a series of 8-bit images in the multi-tiff format using Deeptrack 2.1 [5]. Noise is added to the synthetic images to account for background fluorescence and shot noise. A map corresponding to the segmentation of VIP particles at the first frame for which CPs and diffusion parameters must be detected is also provided as a tiff file. Connected components of the map are labeled with unique integer values that correspond to the particle index.
For Track 2 (3(a)), we provide a csv file for each FOV containing a table whose columns contain trajectory index, time step, x-coordinate, and y-coordinate. Coordinates of simulated trajectories are corrupted with Gaussian noise corresponding to finite (subpixel) localization precision. The trajectories have a maximum length of 200 frames.
Exemplary data for all the models are shown in 4. Files in different tracks labeled with the same experiment and the FOV index (e.g., Track_1/EXP_4/FOV_3.tiff and Track_2/EXP_4/FOV_3.csv) include simulations obtained with the same set of dynamics parameters but do not correspond to the motion of the same set of particles.
.1 Scoring and evaluation
@@
The performance of the methods will be evaluated using specific metrics for each Task. For ranking purposes in the challenge, composite metrics will be used, as described below.
.1.1 Ensemble Tasks
@@
Participation in an ensemble task requires predictions of the type of model used for simulating each experiment, the number of states $S$ of the model, and the parameters of each state. The type of model will be simply evaluated as correct or wrong. The prediction of the number of states will be assessed by measuring the difference with the ground truth. For both the generalized diffusion coefficient and the anomalous diffusion exponent, predictions must include the mean, the standard deviation, and the relative weight of each state. From these values, we will compute the associated multi-modal distributions $P_{\alpha}$ and $P_{K}$. The similarity of these distributions to the ground-truth distributions $Q_{\alpha}$ and $Q_{K}$ will be assessed by means of the first Wasserstein distance ($W_{1}$),
[…]
@@
where $\mathrm{CDF}_{Q}$ is the cumulative distribution function of the distribution $Q$ and $\mathrm{supp}(Q)$ is the support ($\alpha\in(0,2)$ and $K\in[10^{-12},10^{6}]\,\ {\rm pixel^{2}\cdot frame^{-\alpha}}$).
.1.2 Single-trajectory Tasks
@@
Participation in a single-trajectory task requires predictions of the $M$ CPs and the dynamic properties, i.e., the generalized diffusion coefficient $K$, the anomalous exponent $\alpha$, and diffusive-type identifiers of the resulting $M$$+1$ segments. Different metrics will be used to evaluate the methods’ performance.
@@
Following Ref. [51], given a ground-truth CP at locations $t_{({\rm GT}),i}$ and a predicted CP at locations $t_{({\rm P}),j}$, we define the gated absolute distance:
@@
where $\varepsilon_{\mathrm{CP}}$ is used as a fixed maximum penalty for CPs located more than $\varepsilon_{\mathrm{CP}}$ apart. For a set of $M$${}_{\rm GT}$ ground-truth CPs and $M$${}_{\rm P}$ predicted CPs, we can solve a rectangular assignment problem using the Hungarian algorithm [86] by minimizing the sum of distances between paired CPs:
@@
The distance $d_{\rm CP}$ allows to define a pairing metrics:
@@
where $\overline{d_{\rm CP}}$ is the distance associated with having all unassigned predicted CPs at a distance larger than $\varepsilon_{\mathrm{CP}}$ from all ground-truth CPs. This metric measures the presence of spurious CPs and is bound in $[0,\alpha_{\rm CP}]$, taking value $\alpha_{\rm CP}$ if no spurious CPs are present. We also calculate the number of true positive (TP), i.e., the paired true and predicted CPs with a distance smaller than $\varepsilon_{\mathrm{CP}}$. Spurious predictions, i.e., not associated with any ground truth or having a distance larger than $\varepsilon_{\mathrm{CP}}$ are counted as false positive (FP). Ground truth CPs not having an associated prediction at a distance shorter than $\varepsilon_{\mathrm{CP}}$ are considered false negative (FN). Given an experiment containing $N$ trajectories, we compute the overall number of TP, FP, and FN. We then use these values to calculate the Jaccard similarity coefficient (JSC) over the whole experiment as:
@@
For the predicted CPs classified as TP, we also compute the root mean square error (RMSE), defined as:
@@
For the evaluation of methods performance for the estimation of the dynamic properties, we first follow a procedure similar to the one described above for the pairing of the CPs. Predicted CPs are used to define the predicted trajectory segments. We define a distance between predicted and ground-truth segments based on the JSC calculated with respect to their temporal support, where time points at which predicted and ground-truth segments overlap are considered as TP, predicted time points not corresponding to the ground truth as FP, and ground-truth time points not predicted as FN. The Hungarian algorithm is used to pair segments by maximizing the sum of the JSC. Only paired segments are used to calculate metrics assessing methods performance for the estimation of dynamics properties. For the generalized diffusion coefficient $K$, we use the mean squared logarithmic error (MSLE) defined as:
@@
For the anomalous diffusion exponents $\alpha$, we use the mean absolute error (MAE):
@@
[…]
@@
[…]
.2 Metrics for challenge ranking
For ranking purposes, we will use the mean reciprocal rank (MRR) as a summary statistic for the overall evaluation of software performance [42]:

### v3

@@ 4.2 Competition design
  • FOV — A region of the sample where the recording takes place. Particles within the same field of view (FOV) can undergo interactions among themselves and/or with the environment.
  • Video (Video Track only) — Videos corresponding to each FOV.
  • Trajectory (Trajectory Track only) — Trajectory corresponding to the motion of an individual particle.
@@
For both tracks, all particles used in the simulations and located in the FOV are provided/visualized (i.e., full labeling conditions). The effect of blinking or photobleaching was not taken into account.
@@
In each track, participants could compete in two different tasks, as typically done in the analysis of experimental data:
@@
  • Ensemble Task — Ensemble-level predictions providing, for each experimental condition, the model used to simulate the experiment, the number of states, and the fraction of time spent in each state. For each identified state, participants had to determine the mean and standard deviation of the distribution of the generalized diffusion coefficients $K$, and the mean and standard deviation of the distribution of the anomalous diffusion exponent $\alpha$ corresponding to the underlying motion.
@@
  • Single-trajectory Task — Trajectory-level predictions providing for each trajectory a list of $M$ inner CPs delimiting $M+1$ segments with different dynamic behavior. For each segment, participants had to identify the generalized diffusion coefficient $K$, the anomalous diffusion exponent $\alpha$ corresponding to the underlying motion, and an identifier of the kind of constraint imposed by the environment ($0=$ immobile, $1=$ confined, $2=$ free (unconstrained, $0.05\leq\alpha<1.9$), $3=$ directed ($1.9\leq\alpha<2.0$). For the Video Track, predictions had to be provided for a subset of particles (in the following, we will refer to them as VIP, very important particles) identified through a label map of the first frame of the movie. For the Trajectory Track, predictions had to be provided for all trajectories in the FOV.
@@
For each task, several metrics were evaluated (see Scoring and evaluation). Participants were allowed to provide partial submissions, e.g., including predictions for a limited subset of experiments or for specific parameters. For ranking purposes of the Challenge, missing predictions were scored with the worst possible value of the corresponding metric.
4.3 Competition overview
The 2nd AnDi Challenge was held between December 1, 2023 and July 15, 2024 on the Codalab platform. It was divided into three phases, namely Development, Validation, and Challenge. The Development Phase (5 months) was intended for the participants to set up their methods, test them, and familiarize themselves with the datasets and the scoring platform. An unlabeled dataset was available and the public leaderboard showed scores obtained on this dataset. An online workshop was held on February 22, 2024 to instruct the participants about the details of the challenge. The Validation Phase (1 month) was a test of the actual final challenge. A new dataset (described in Challenge Dataset) was provided and the leaderboard was again public. The Challenge Phase (15 days) was the final stage of the competition. A new dataset was provided and the number of submissions per team was limited to 1 per day. The results were not publicly disclosed and the leaderboard was made public only after the end of the competition. In total, we received 1343 submissions during the three phases. Participants registered in teams of 1 to 5 people. In the final stage, out of 80 registered participants, 53 individuals, divided into 18 teams, were included in the leaderboard (see Supp. Tab. 2 for the list of participating teams). The teams’ affiliations spanned Europe (12 teams), Asia (6 teams), and America (1 team). From the final leaderboard, members of the top 5 teams in each task were invited to co-author this article. An overview of these teams and the methods is provided in Supp. Infor. - Overview of Teams and Methods.
@@ 4.3 Competition overview
The results of the Challenge were discussed with the participants and other experts from the field during the $2^{\rm nd}$ Anomalous Diffusion Workshop that was held in June 2025.
4.4 Challenge Dataset
The Challenge dataset was composed of 12 experiments corresponding to different diffusion models and parameter values. Details about the numeric values of parameters of the experiments are given in Supp. Tab. 3. In addition, Supp. Fig. 1 summarizes the distribution of specific features within the dataset. EXP 1 aimed at mimicking multistate diffusion in membrane proteins. Average diffusion coefficients and the transition matrix of the MSM were chosen to reproduce, with the appropriate scaling, the three fastest states reported for the diffusion of the $\alpha$2A-adrenergic receptor [80]. EXP 2 reproduced changes in diffusion coefficient due to protein dimerization, inspired by the behavior reported for the epidermal growth factor receptor ErbB1 [77]. EXP 3, EXP 4, and EXP 5 were designed to compare the methods’ ability to detect changes from the same free diffusive state to a slow diffusing state characterized either by traps (QTM, EXP 3), small confinement regions (TCM, EXP 4), or a subdiffusive dimeric state (DIM, EXP 5). EXP 6 and EXP 7 were meant to assess the methods’ ability to take advantage of the knowledge of the physical model itself and additional information present in the experiment to improve predictions. The experiments corresponded to different theoretical models (DIM and MSM) with the same diffusive parameters. EXP 8 served as a negative control and contained only SSM trajectories with very broad distributions of $K$ and $\alpha$. EXP 9 was generated from QTM with very short trapping times and superdiffusion in the free state to assess how the methods deal with such extreme conditions. The other three experiments contained data with extreme and unrealistic parameters meant to assess potential biases of the methods, and will not be discussed further.
4.5 Scoring and evaluation
@@
The performance of the methods was evaluated using specific metrics for each task. For ranking purposes in the Challenge, composite metrics were used, as described below.
4.5.1 Ensemble Task
@@
Participation in the Ensemble Task required predictions of the type of model used for simulating each experiment, the number of states $S$ of the model, and the parameters of each state. The type of model was simply evaluated as correct or wrong. The prediction of the number of states was assessed by measuring the difference with the ground truth. For both the generalized diffusion coefficient and the anomalous diffusion exponent, predictions had to include the mean, the standard deviation, and the relative weight of each state. From these values, we computed the associated multi-modal distributions $P_{\alpha}$ and $P_{D}$. The similarity of these distributions to the ground-truth distributions $Q_{\alpha}$ and $Q_{D}$ was assessed by means of the first Wasserstein distance ($W_{1}$),
$W_{1}(P,Q)=\int_{\mathrm{supp}(Q)}|\mathrm{CDF}_{P}(x)-\mathrm{CDF}_{Q}(x)|dx$ (3)
@@
where $\mathrm{CDF}_{Q}$ is the cumulative distribution function of the distribution $Q$ and $\mathrm{supp}(Q)$ is the support ($\alpha\in(0,2)$ and $K\in[10^{-12},10^{6}]\,\ {\rm pixel^{2}/frame^{\alpha}}$).
4.5.2 Single-trajectory Task
@@
Participation in the Single-trajectory Task required predictions of the $M$ CPs and the dynamic properties, i.e., the generalized diffusion coefficient $K$, the anomalous exponent $\alpha$, and diffusive-type identifiers of the resulting $M+1$ segments. Different metrics were used to evaluate the methods’ performance.
@@
Following Ref. [51], given a ground-truth CP at locations $t_{({\rm GT}),i}$ and a predicted CP at locations $t_{({\rm P}),j}$, we defined the gated absolute distance:
@@
where $\varepsilon_{\mathrm{CP}}$ was used as a fixed maximum penalty for CPs located more than $\varepsilon_{\mathrm{CP}}$ apart. For a set of $M_{\rm GT}$ ground-truth CPs and $M_{\rm P}$ predicted CPs, we solved a rectangular assignment problem using the Hungarian algorithm [85] by minimizing the sum of distances between paired CPs:
@@
The distance $d_{\rm CP}$ allows to define a pairing metric:
@@
where $\overline{d_{\rm CP}}$ is the distance associated with having all unassigned predicted CPs at a distance larger than $\varepsilon_{\mathrm{CP}}$ from all ground-truth CPs. This metric measures the presence of spurious CPs and is bound in $0,\alpha_{\rm CP}]$, taking value $\alpha_{\rm CP}$ if no spurious CPs are present. We also calculate the number of true positives (TP), i.e., the paired true and predicted CPs with a distance smaller than $\varepsilon_{\mathrm{CP}}$. Spurious predictions, i.e., not associated with any ground truth or having a distance larger than $\varepsilon_{\mathrm{CP}}$ were counted as false positives (FP). Ground truth CPs not having an associated prediction at a distance shorter than $\varepsilon_{\mathrm{CP}}$ were considered false negatives (FN). Given an experiment containing $N$ trajectories, we computed the overall number of TP, FP, and FN. We then used these values to calculate the JSC over the whole experiment as:
@@
For the predicted CPs classified as TP, we also computed the root mean square error (RMSE), defined as:
@@
For the evaluation of the methods’ performances on the estimation of the dynamic properties, we first followed a procedure similar to the one described above for the pairing of the CPs. Predicted CPs were used to define the predicted trajectory segments. We defined a distance between predicted and ground-truth segments based on the JSC calculated with respect to their temporal support, where time points at which predicted and ground-truth segments overlap were considered as TP, predicted time points not corresponding to the ground truth as FP, and ground-truth time points not predicted as FN. The Hungarian algorithm was used to pair segments by maximizing the sum of the JSC. Only paired segments were used to calculate metrics assessing methods performance for the estimation of dynamics properties. For the generalized diffusion coefficient $K$, we used the mean squared logarithmic error (MSLE) defined as:
@@
For the anomalous diffusion exponents $\alpha$, we used the mean absolute error (MAE):
@@
where $N$ is the total number of paired segments in the experiment, $\alpha_{{\rm(GT)},i}$ and $\alpha_{{\rm(P)},j}$ represent the ground-truth and predicted values of the anomalous exponent of paired segments, respectively. For the classification of the type of diffusion, we used the F1-score:
@@
where TPc, FPc, and FNc represent true positives, false positives, and false negatives with respect to segment classification. The metric was calculated as a micro-average, which aggregates the contributions of all classes to compute the average metric and is generally preferable when class imbalance is present.
4.6 Metrics for challenge ranking
For ranking purposes, we used the mean reciprocal rank (MRR) as a summary statistic for the overall evaluation of software performance [42]: